Nittany Company uses a periodic inventory system. At the end of the annual accounting period, December 31 of the current year, the accounting records provided the following information for product 1: 

                                                 Units  Cost  

Inventory, December 31, prior year 2,000 Unit $3

For the current year: 

Purchase, March 21                     5,190  5

Purchase, August 1                      2,910  6

 Inventory, December 31, current year   4,040 

Required: Compute ending inventory and cost of goods sold for the current year under FIFO, LIFO, and average cost inventory costing methods. (Round "Average cost per unit" to 2 decimal places and final answers to nearest whole dollar amount.) 
Answer #1

FIFO periodic inventory system:

1. closing stock Included latest purchases

2. cost of goods sold is Included oldest purchases

Ending inventory value = Ending units inventory x Latest purchase Rate (August 1)

= 4040 X 6

=24240

Cost of goods sold

= Opening Inventory + Value of purchases – Closing Inventory

= (2000 X 3 + 5190 X 5 + 2910 X 6 - 24240)

=6000 + 25950 + 17460 - 24240

= 25170

LIFO periodic inventory system, closing stock included oldest purchases and cost of goods sold is included latest purchases

Ending inventory value = Ending units inventory x oldest inventory Rate

= (2000 X 3 + 2040 X 5)

= 6000+ 10200 = 16200

Cost of goods sold

= Opening Inventory + Value of purchases – Closing Inventory

= (2000 X 3 + 5190 X 5 + 2910 X 6 - 24240)

=6000 + 25950 + 17460 - 16200

= 33210

Average cost per unit

= Total value of Inventory / Total Units =(6000 + 25950 + 17460 ) /(2000+5190+2910)

= 49410/10100

=4.90 PER UNIT

value of ending inventory

Ending inventory value = Ending units inventory x Cost per unit

= 4040 X 4.90 = 19796

Cost of goods sold

= Opening Inventory + Value of purchases – Closing Inventory

= (2000 X 3 + 5190 X 5 + 2910 X 6 - 19796

= 29614